[figure: 9.7 cm]

THE ROLE OF ORBITAL ON THE MAGNETIC PROPERTIES

OF SOME V OXIDES

G.A. Sawatzky

Laboratory of Applied and Solid State Physics, Materials Science Center, University of Groningen, Groningen, The Netherlands

Orbital degeneracy plays a very important role in the magnetic properties of transition metal compounds. In recent years a lot of attention has been place on the relation between the magnetic structures and the orbital occupation of the 3d transition metal compounds. Resonant X ray scattering has made it possible to quite directly determine the local distortions related to orbital ordering and recent studies of V_{2}O_{3} , YVO_{3} have revealed some surprises. Also recent studies of the properties of LiVO_{2} and YVO_{3} display surprising temperature dependent behaviour such as transitions to non magnetic states and temperature dependent multiple magnetisation reversals. I will discuss these results and present an interpretation in terms of the strong coupling between the orbital and spin structure. Recent references are : Y. Ren et al., Nature 396, 441 (1998) H.F. Pen et al., PRL 78, 1323 (1997) I.S. Elfimov et al., PRL June 1999 M.A. Korotin et al., PRL 80, 4305, (1998)

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : PLENARY LECTURE

Corresponding author : G.A. Sawatzky

Address for correspondence : Laboratory of Applied and Solid State Physics, Materials Science Center, University of Groningen, Groningen, The Netherlands

Email address : G.A.Sawatzky@phys.rug.nl